OpenStudy (anonymous):

find the point (x,y) on the graph of the equation y=2√x which is closest to the point (11,0). use this to write down the minimum distance from the graph to the point (11,0.)

OpenStudy (dumbcow):

use distance equation \[d^2 = (x-11)^2 + (2 \sqrt{x} - 0)^2\] \[d^2 = x^2 -18x + 121\] differentiate \[2d d' = 2x - 18\] \[d' = \frac{2x - 18}{2d} = \frac{x-9}{d}\] minimize distance by setting derivative equal to zero \[\frac{x-9}{d} = 0 \rightarrow x = 9\] point closest to (11,0) is at x=9 \[(9,6)\] min distance is: \[\sqrt{40} = 2\sqrt{10}\]